# Rethinking the Relationship between the Power Law and Hierarchical Structures

### v1

@@ 5 Discussion and Conclusion
Lin and Tegmark (2017) attributed the power-law decay of correlation in natural languages to underlying hierarchical structures, which has inspired numerous studies on both human languages and animal signals. However, their explanation raises several important questions, as discussed in Section 2.3. To address these, the present study analyzes the statistical properties of hierarchical syntactic structures in natural languages, to test the questions essential for the proposed argument.
@@
The analyses answer all the questions negatively: (i) The correlation in syntactic structures decays according to a power law with respect to the structural distance; (ii) The growth of the sequential distance is slower than linear with the structural distance, due to the strong branching bias in syntactic structures; (iii) The statistical properties of syntactic structures differ from those of PCFGs, particularly in the CFIB. These results reveal that the argument by Lin and Tegmark (2017) cannot account for the mechanism underlying the power law at the sentence level.
@@
We also analyze the PCFG that approximates the parse trees. The behavior of CFIB in the PCFG clearly differs from that in the parse trees. Furthermore, this analysis demonstrates that the asymptotic behavior of the PCFG does not emerge for short sequences. This result highlights the difficulty of applying the proposed argument not only to single sentences in adult human languages but also to child languages and animal signals.
While further investigation is needed for other parsing schemes and corpora, we expect qualitatively similar results in more general settings. Indeed, we have confirmed similar results for unbinarized parse trees and for the WikiText corpus (Merity et al., 2017), see Appendices C and D for details. Whether these behaviors appear in other languages remains an open question. A key challenge is the limited availability of treebanks for languages other than English.
Another important open question is whether the argument by Lin and Tegmark (2017) applies at scales longer than sentences, that is, at the level of entire texts, where the statistical behaviors of PCFGs may be sufficiently close to their asymptotic behaviors. To extend our analysis to the level, a necessary first step is to formalize how to describe the hierarchical structures of entire texts and develop datasets annotated with such information.
@@
Since the discussion in previous research does not account for the power law in natural languages, at least at the sentence level, alternative explanations are also needed. One promising direction is critical phenomena in statistical physics. These refer to the phenomena that statistical quantities exhibit power-law behaviors near phase transition point, where statistical properties of a system change qualitatively. Indeed, as the temperature parameter of large language models varies, generated texts exhibit critical phenomena, sharing statistical properties with natural languages, including the power-law decay of correlation (Nakaishi et al., 2024).
Finally, more detailed characterizations of the statistical quantities are necessary. This study used simple fitting functions to distinguish between exponential and power-law behaviors. However, the observed statistical quantities exhibit more complex patterns. A closer examination of these patterns will be essential for understanding syntactic structures through statistical analysis.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Nos. 23H01095, 23KJ0622, and 24H00087; JST, PRESTO Grant No. JPMJPR21C2; and the World-Leading Innovative Graduate Study Program for Advanced Basic Science Course at the University of Tokyo.
Appendix A Tag Reduction for Parse Trees
To reduce the number of node states, we grouped tags with similar functions into the same category. The eleven categories used in our analyses are shown in Table 1.
@@ Appendix A Tag Reduction for Parse Trees
Table 1: Eleven tags used in this study and the corresponding part-of-speech tags and phrase labels.
Appendix B Context-free Independence Breaking
[…]
Appendix C Unbinarized Case
As discussed in Section 3.1, we binarized syntactic trees during preprocessing. To verify that this binarization does not significantly affect our results, we analyze the statistical properties of the unbinarized trees in this section. The MI in sequences is unaffected by binarization. In addition, the dataset is too limited to reliably compute the CFIB for unbinarized trees. Therefore, we focus on the MI in structures and the growth of sequential distance. The results, shown in Figure 10, exhibit the power-law decay of the MI and the slow growth of the sequential distance, which are qualitatively consistent with those observed in the binarized case, although, as expected, the structural distances are smaller. These suggest that the statistical behavior of syntactic structures is robust to binarization.
Figure 9: CFIB $J$ for (a) $(x_{0},x_{1})=(\text{NP},\text{VP})$ and (b) $(x_{0},x_{1})=(\text{VP},\text{VP})$. Fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$ in (a) and $N_{\text{data}}=6.4\times 10^{5}$ in (b).
Figure 10: Results for unbinarized parse trees. (a) MI $I_{\text{tag}}$ between POS or phrasal tags. Fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$. (b) Distribution and average of sequential distances.
@@ Appendix C Unbinarized Case
Figure 11: Results for WikiText. (a) MI $I_{\text{POS}}$ between POS tags. (b) MI $I_{\text{tag}}$ between POS or phrasal tags. (c) Distribution and average of sequential distances. (d) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$. Fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$ in (a), (b), and (d).
Appendix D WikiText
In the main text, we analyzed the BLLIP corpus. To address whether similar behaviors are observed in other corpora, we perform the same analyses for the WikiText-103 corpus (wikitext-103-raw-v1 split) (Merity et al., 2017), which is a collection of English Wikipedia articles that preserves the original Wikipedia markup and contains over 100 million tokens and approximately 3.9 million sentences. We implement a rule-based detokenizer and segment sentences with a spaCy sentencizer (Honnibal et al., 2020). Then, we extract sentences up to 40 words (resulting in approximately 3.4 million sentences) and apply the Berkeley Neural Parser (benepar_en3 model) (Kitaev and Klein, 2018; Kitaev et al., 2019) to them to obtain Penn Treebank-style constituency trees. Finally, we binarize the trees and reduce the number of node states using the same method applied to the BLLIP corpus.
Figure 11 shows that the results are qualitatively similar to those for BLLIP: the MI decays according to a power-law function in both sequences and structures; the growth of sequential distance is slower than linear; and the CFIB also decays following a power law. These suggest that the observed behaviors are universal, and that the assumption in Lin and Tegmark (2017) does not hold for syntactic structures across corpora.
However, we note a slight difference in the behavior of MI in both sequences and structures: it increases at large distances, $r_{\text{seq}}\gtrsim 30$ and $r_{\text{str}}\gtrsim 20$. This may be due to structural differences in unusually long sentences or potential parsing errors in such cases. We are unable to determine the cause of this effect, as the dataset contains too few long sentences.

### v2

@@ 5 Discussion and Conclusion
In natural languages, the correlation decays as a power law with respect to the sequential distance, as reported in numerous previous studies. At the same time, linguists generally agree that natural languages exhibit hierarchical syntactic structures. Although both phenomena are well established, their relationship remains elusive.
@@
The present study examines this relationship, specifically the argument proposed by Lin and Tegmark (2017). They attributed the power-law decay of correlation in natural languages to underlying hierarchical structures, which has since inspired numerous studies on both human languages and animal signals. However, their argument relies on several assumptions that are questionable from a linguistic perspective. To examine the validity of the assumptions for syntactic structures, we test the three key questions through statistical analyses of parse trees in natural languages.
@@
The analyses yield negative answers to all three questions: (i) The correlation in syntactic structures decays according to a power law with respect to the structural distance; (ii) The growth of the sequential distance is slower than linear with the structural distance, due to the strong branching bias in syntactic structures; (iii) The statistical properties of syntactic structures differ from those of PCFGs, particularly in the CFIB. These results reveal that the relationship between the power-law decay of correlation and syntactic structures differs from that proposed by Lin and Tegmark (2017).
We obtained similar results under various settings: for the English treebank BLLIP with several different schemes, for another English corpus, WikiText, and for the Japanese treebank NPCMJ. The consistency of the results across different schemes, domains, and languages suggests the robustness of our conclusion.
@@
We also analyze the PCFG obtained from the maximum likelihood estimation of BLLIP. The statistical properties of the PCFG clearly differ from those in parse trees. Furthermore, this analysis demonstrates that the exponential decay of $I_{\text{tag}}$ in the PCFG does not emerge within individual sentences. Although Lin and Tegmark (2017) proved that $I_{\text{tag}}$ decays exponentially at sufficiently large structural distances, the structural distances in typical parse trees are too short for this behavior to appear. This result highlights the difficulty of applying the proposed argument not only to sentences produced by human adults but also to child speeches and animal signals, in which the sequences are shorter than sentences by human adults.
In addition, the statistical differences between natural language parse trees and trees generated by PCFGs may have implications for theoretical research on large language models. Some studies in this field analyze models trained on synthetic data, in which PCFGs are sometimes used to generate the data (Allen-Zhu and Li, 2023; Cagnetta and Wyart, 2024). However, the model’s ability to learn PCFGs will not account for its ability to learn syntactic structures of natural languages, although some studies employ PCFGs as models of more abstract structures than syntax.
Currently, we cannot rule out the possibility that the MI in natural language syntactic structures follows an exponential decay at structural distances much larger than those observed in this study. However, even if this is the case, such large parse trees would be too rare to have a significant influence on the statistical properties of the corpus as a whole. Moreover, as long as syntactic structures are biased, the decay of MI in sequences should be faster than a power law, because the sequential distance grows more slowly than exponentially. Therefore, statistical analyses of larger parse trees are unlikely to reveal the mechanism underlying the power-law decay of correlation in natural languages.
A more promising direction is to consider structures larger than individual sentences, namely, the hierarchical organization of entire texts composed of multiple sentences. The argument by Lin and Tegmark (2017) may account for the power-law decay of correlation observed across entire texts rather than within single sentences. As a first step in this direction, it will be necessary to develop sufficiently large-scale corpora annotated with discourse information to estimate statistical quantities such as $I_{\text{POS}}$, $I_{\text{tag}}$, and $J$.
@@
Another direction is to explore approaches different from that of Lin and Tegmark (2017). One alternative is critical phenomena in statistical physics. These refer to the phenomena that statistical quantities exhibit power-law behaviors near phase transition points, at which statistical properties of a system change qualitatively. Indeed, as the temperature parameter of large language models varies, generated texts exhibit critical phenomena, sharing statistical properties with natural languages, including the power-law decay of correlation (Nakaishi et al., 2024).
@@ Acknowledgments
We thank Taiga Ishii, Shinnosuke Isono, Sho Yokoi, and Keisuke Sakaguchi for useful discussions and suggestions. This work was supported by JSPS KAKENHI Grant Nos. 23H01095, 23KJ0622, 24H00087, and 25K24434; JST, PRESTO Grant No. JPMJPR21C2; and the World-Leading Innovative Graduate Study Program for Advanced Basic Science Course at the University of Tokyo.
Appendix A Tag Reduction
To reduce the number of categories, we grouped tags with similar functions into the same category. The eleven categories used for BLLIP and WikiText and the seven categories for NPCMJ are shown in Tables 1 and 2, respectively.
@@ Appendix A Tag Reduction
Table 1: Eleven tags used for BLLIP and WikiText and the corresponding part-of-speech tags and phrase labels.
Tags Part-of-Speech Tags and Phrase Labels
NP NP NUMCLP PRN NML PNLP NUMCLPSYM N NPR NUM CL ADJN PRO ADJI D Q FN WPRO PNL PRN WNUM WD FW
IP IP VB AX AXD VB0 VB2 PASS PASS2 MD
$P$ PP P
ADVP ADVP ADV NEG WADV
CONJP CONJP CONJ
$P$ PP P
CP CP FRAG INTJP FS INTJ
Others LST META LS PU PUL PUR SYM PUQ COMMENT
Table 2: Seven tags used for NPCMJ and the corresponding part-of-speech tags and phrase labels.
Appendix B Other Schemes
In the main text, we analyzed parse trees after applying two preprocessing steps. First, we binarized the parse trees. Second, we reduced the number of tag types by grouping similar tags, where we use the same reduced categories for both POS and phrasal tags.
Although the statistical quantities depend on the specific preprocessing scheme, we expect that these details do not affect our conclusions, such as whether the MI decays according to a power law or an exponential function, and whether the CFIB is significantly positive. To examine this expectation, we perform additional analyses under two alternative preprocessing schemes: the unbinarized setting and the phrasal-tag-only setting, in which POS and phrasal tags are distinguished and only the latter are analyzed.
@@ Appendix B Other Schemes
Figure 11: (a1, 2) Results for the unbinarized setting: (a1) MI $I_{\text{tag}}$ between POS or phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) distribution and average of sequential distances. (b1, 2) Results for the phrasal-tag-only setting: (b1) MI $I_{\text{tag}}$ between phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) CFIB $J$, where fitting was pwerformed for $N_{\text{data}}=6.4\times 10^{5}$.
Figures 11 (a1) and (a2) show the results for the unbinarized setting. Because binarization does not affect the MI in sequences and the dataset is too limited to reliably estimate the CFIB for unbinarized trees, we present only the MI in structures and the sequential distance as a function of the structural distance. For both quantities, the results are qualitatively similar to those for the binarized setting in the main text: the decay of $I_{\text{tag}}$ follows a power-law function rather than an exponential function, and the growth of $r_{\text{seq}}$ is slower than exponential.
In the phrasal-tag-only setting, the MI between POS tags and the sequential distance are not defined. We therefore estimate the MI between phrasal tags and the CFIB, presented in Figures 11 (b1) and (b2), respectively. Both quantities exhibit power-law decay, consistent with the results in the main text.
Across both alternative preprocessing schemes, our main conclusion holds: the assumptions in the argument by Lin and Tegmark (2017) are not satisfied. These results indicate that our findings are robust to the details of preprocessing.
Appendix C Alternative Estimators
In the main text, we computed the MI using the estimator of Grassberger (2003) (GR), whose bias decreases rapidly. The estimated values saturate as $N_{\text{data}}$ increases, except for the CFIB in the PCFG. The estimates of the CFIB in the PCFG continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, which is reasonable because the true value is zero. These results show that we have properly estimated the statistical quantities.
To further support the reliability of our estimates, we compute $I_{\text{POS}}$, $I_{\text{tag}}$, and $J$ using six alternative estimators proposed by Chao and Shen (2003) (CS), Horvitz and Thompson (1952) (HT), Miller (1955) and Madow (1948) (MM), Nemenman et al. (2001) (NSB), Wolpert and Wolf (1995) (WW) and Zahl (1977) (ZA). For these computations, we use the Python package entropy-estimation (Arora et al., 2022).
Figure 12 presents the results for BLLIP computed with the six alternative estimators, together with those obtained using the GR estimator, which are the same as in the main text. The estimates agree with ours in most cases. Figures 12 (b) and (c) show that the WW estimates for $I_{\text{tag}}$ and $J$ at large distances exhibit clear bias, producing some negative estimates even though the true MI and CFIB are always non-negative. The CS and HT estimates for $J$ also exhibit similar behaviors. These results are consistent with Arora et al. (2022), which reports that these three estimators yield larger errors than the others when the sample size is large. From these results, we can confirm that our estimates in the main text are reliable.
Figure 12: Results for BLLIP obtained with different estimators. (a) MI $I_{\text{POS}}$ between POS tags. (b) MI $I_{\text{tag}}$ between POS or phrasal tags. (c) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$. In all cases, $N_{\text{data}}=2.56\times 10^{6}$.
Figure 13: Results for the PCFG obtained with different estimators. (a) MI $I_{\text{POS}}$ between POS tags. (b) MI $I_{\text{tag}}$ between POS or phrasal tags. In both (a) and (b), $N_{\text{data}}=8\times 10^{6}$. (c-e) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$ obtained with (c) $M$, and (d) NSB, and (e) ZA.
Figure 13 shows the results for the same quantities for the PCFG obtained from the maximum likelihood estimation of BLLIP. The estimates of $I_{\text{POS}}$ and $I_{\text{tag}}$, shown in Figures 13 (a) and (b), are consistent across all estimators. This again supports the reliability of the results in the main text.
We also show the estimates of the CFIB $J$ obtained with the MM, NSB, and ZA estimators in Figures 13 (c), (d), and (e), respectively. The MM and ZA estimates continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, similar to the GR estimates in the main text. This means that these two estimates properly converge to zero, the true value of the CFIB in PCFGs. The convergence of the NSB estimates is rather unclear. We do not present results with CS, HT, and WW, because most of their estimates are negative and cannot be plotted on a log scale.
In summary, MM, ZA, and GR, which is used in the main text, perform well for all statistical quantities. NSB generally performs well, yet its convergence to the true value is unclear for the CFIB in the PCFG. The other estimators, CS, HT, and WW, sometimes exhibit clear bias, implying that they should be avoided when the sample size is large.
Appendix D Context-free Independence Breaking
In Section 4.4, we presented the results of CFIB for BLLIP when the two nodes are fixed as $(x_{0},x_{1})=(\text{NP},\text{NP})$. In this section, we present the results with $(x_{0},x_{1})=(\text{NP},\text{VP})$ and $(\text{VP},\text{VP})$ to examine how the metric depends on the fixed nodes.
From the results shown in Figure 14, we can observe a power-law decay similar to that with $(x_{0},x_{1})=(\text{NP},\text{NP})$, although the power law exponents may be different. This implies that the CFIB in syntactic structures follows a power-law decay, irrespective of the fixed nodes.
Figure 14: CFIB $J$ for BLLIP with (a) $(x_{0},x_{1})=(\text{NP},\text{VP})$ and (b) $(x_{0},x_{1})=(\text{VP},\text{VP})$. Fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$ in (a) and $N_{\text{data}}=6.4\times 10^{5}$ in (b).